4He beam tuning for 10Bap experiment (Nov. 9 2020)

Nov 9, 2020 • pelletron

14:00

  • Turn on RF source after fixing extractor entrance holder and gas bottle valve not open.
  • Beam current stable at PRB VC=4.5 kV and EXT VC =5 kV but very low ~ 1.74 nA.

14:58

  • Try increasing PRB VC to 6 kV and EXT VC to 5.8 kV -> Stable current at 12 nA : still too low but increasing.
  • gas pressure at IGC 01-1 at 1.6e-6 T
  • IM CC set at 7.03 A.
  • See attachment 1 for beam condition

15:50

  • Try decrease gas pressure at IGC 01-1 to 1.4e-6 T..Beam current at FC01-1 now ~12.6 nA

15:56

  • Try increase OVN temperature. Now set OVN VC at 54 V.
  • Beam current at FC01-1 suddenly increased t 200 nV and then decrease (see attachment 2). Now at ~60 nA.

14:06

  • Beam continue to decrease
  • Try increasing temperature. Now set OVN VC at 57 V.
  • Beam condition: see attachment 3

16:35

  • Beam condition see attachment 4: beam current at FC01-1 average at 38 nA, quite unstable but Ok for now.

16:45

  • PRB current suddenly drop and beam current at FC01-1 becomes more stable.at 31 nA (see attachment 5)

16:47

  • Beam current at FC01-1 is now about 38 nA.
  • RF source shutdown

elog1 elog2